Dive Number: 3946
December 5, 2003
Port Observer: Tim Shank
Starboard Observer: Gavin Eppard
The purpose of this dive was to survey the deployed ASU unit at Ty and Io and if well developed, collect unit with chemistry, conduct chemistry on the scar, deploy new unit in same place, run plume geochemical survey at Ty and Io, collect Fries #22 at just south of the Bio9 smoker complex, transit to TICA and collect 3 baby traps, collect mussels and clams as possible, transit to the LVWS elevator 379 m to the west, and deploy at TICA.
This dive was extremely interesting given the changes in the biological community structure and black smokers that have taken place in this region. The chimneys that scientists have been calling “Ty" and "Io” are actually Io and Alvinellid Pillar. This is confirmed given the presence of French marker 15 we observed at the base of the Alvinellid Pillar. There have been dramatic changes in the Biomarker 82 region of the BioTransect — mussels now completely dominate this area; no Riftia were observed on the floor of the ASCT. We surveyed the Alvinellid pillar frying pan and the one at Io. We recovered the Io frying pan but not the one at the Alvinellid Pillar. We conducted electro-chemistry on the side and top of the ASU IO frying pan and conducted geochem with the Sipper wand as well and took 2 Sipper samples (side and top). We collected the pan and placed in forward biobox, repeated chemistry on the scar, and deployed a new unit in same place. We then conducted a plume geochemical survey outside of vent plume at Io, at the plume seawater interface and directed in the plume at the 33–40°C height. We then transited to P Vent to survey for alvinellids and conduct our phone patch to the classrooms. After the phone call, we moved north and collected Fries #22 (and took video of the Rusty Riftia vemcos and Biomarker 9 Riftia). We then transited to the elevator and set it up at TICA for recovery the following morning.
GMT Time Activity
14:11 Free of ship. 2 ground on the AUX.
15:20 15:29 On the descent, we started recording electrode data survey. Each of the electrodes continued recording data until we reached the bottom. Was not instructed until five days later that running the flow cell electrode stresses it when in DI water. So, this may explain why electrode #2 did not function.
15:39 2504 On bottom. On a terrain of large pillows with a moderate sediment cover (a light dusting on pillow surfaces, but a complete covering in the interstices). Sea cucumbers, crinoids, and short-stalked sponges with oval-shaped, clear gelatinous tops were present on the surface of the rocks and occasional red shrimp were in the seafloor cracks and in the water column. A strong current from north to south was evident. Upper lift prop has a 4 ground.
15:55 E-chem scans in transit outside of the axial summit collapse trough to Io.
16:00 E-chem scans in front of the Alvinellid Pillar.
16:20 17:14 E-chem survey on the side and top of Io vent. Temperatures ranged from 25 to 80°C. Took Hi Definition video through the pilot’s viewport. Collected Sipper sample #1 and #2 on side and top of ASU frying pan, respectively.
17:21 Collected the ASU frying pan at Io vent and placed in forward biobox.
17:28-17:38 E-chem on the scar created by removing the ASU frying pan (119-138°C); collected Sipper samples and digital photographs of the scar. Collected Sipper sample #3.
17:43-17:53 E-chem survey of the Io vent plume. The Macro-Wand measurements were made outside of the buoyant plume (3.5–4°C), on the boundary interface between plume and deep-sea water (6.5–10°C), and in the vent plume (43–50°C) and again on the opposite boundary (7.5–10°C). Took digital photographs.
17:57 Deployed new ASU frying pan at Io vent and took 3-chip video.
18:02 18:15 Conducted e-chem survey around the new frying pan (side and top — 1 cm above unit) between 20°C and 51°C. Took Sipper sample #4 on top of new ASU unit.
18:20 traversed to P vent to survey allvinellid worms and conduct the phone patch to the classrooms.
19:33 leaving P vent for Bio9 Vent to search for Fries #22.
19:48 Collected Fries #22. The “22” marker broke away while in Alvin’s claw, so this was not recovered.
20:13 Traversing to elevator 379 meters away from Bio9.
20:24 At LVWS elevator and transiting to TICA.
20:45 Arrived at TICA to deploy elevator in Riftia community.
20:53 Pulling weights off of the elevator. The pilot had a difficult time situating the hose for water sampling. Finally found ~7.5°C water and placed the suction hose in this fluid.
21:28 LVWS deployed with the switch placed in the “on” position.
21:30 Weights away.
|Copyright University of Delaware, November 2003|